# Reflected generalized BDSDEs driven by non-homogeneous Lévy processes and obstacle problems for stochastic integro-PDEs with nonlinear Neumann boundary conditions

### v2

@@ 4.1 Reflected SDE driven by inhomogeneous Lèvy process
Let $G$ be an open connected bounded domain of $\mathbb{R}^{l}$ $(l\geq 1)$. Let us assume that $G$ is a smooth domain such that for a function $\Psi\in C^{2}_{b}(\mathbb{R}^{l})$, $G$ and $\partial G$ are defined as $G=\left\{x\in\mathbb{R}^{l}:\Psi(x)>0\right\}$ and $\partial G=\left\{x\in\mathbb{R}^{l}:\Psi(x)=0\right\}.$ For any $x\in\partial G$, $\nabla\Psi(x)$ represents the unit normal vector directed towards the interior of $G$. Moreover, the interior sphere condition is satisfied (see [38], pp.551), meaning there exists a $m>0$ such that for each $x\in\partial G$ and $x^{\prime}\in\bar{G}$, the following holds:
@@
For every $(t,x)\in[0,T]\times\bar{G}$, we consider the subsequent reflected stochastic differential equation driven by the inhomogeneous Lèvy process $L$ (SDEL):
@@
where $\sigma:\mathbb{R}^{l}\to\mathbb{R}^{l}$ is globally Lipschitz, i.e. $\exists C>0,\forall x,x^{\prime}\in\mathbb{R}^{l},|\sigma(x)-\sigma(x^{\prime})|\leq C|x-x^{\prime}|,$ and satisfies the linear growth condition: $|\sigma(x)|\leq C(1+|x|)$.
@@
Furthermore, we present the subsequent result regarding the existence and uniqueness of the solution to (4.49). The argument adheres to the same methodology as outlined in [23]. For the reader’s convenience, we delineate the principal steps of the proof.
@@
The uniqueness follows directly from Proposition 14. For existence, we do the following. For the sake of brevity, we shall denote $\mathfrak{R}=\mathfrak{R}^{t,x}$ where $\mathfrak{R}=X,\kappa$. For each $n\geq 0$, define the truncated process:
$L_{t}^{n}=\int_{0}^{t}b_{s}ds+\int_{0}^{t}\int_{\frac{1}{n}<|e|<1}e\mu(ds,de)-\int_{0}^{t}\int_{\frac{1}{n}<|e|<1}eF_{s}(de)ds.$ (4.50)
Given that $L^{n}$ possesses finite variation, the Lipschitz and growth conditions are satisfied on $\sigma$, and the domain $G$ adheres to the interior sphere condition (4.48), it follows from [23] that for each $n$, a unique solution $(X^{n},\kappa^{n})$ exists for the truncated problem:
$X_{s}^{n}=x+\int_{t}^{s}\sigma(X_{r-}^{n})dL_{r}^{n}+\int_{t}^{s}\nabla\Psi(X_{r}^{n})d\kappa_{r}^{n}.$ (4.51)
Furthermore, according to Proposition 14 and Jakubowski’s compactness requirement for processes exhibiting right-continuity with left limits (see [21]), there exists a subsequence $(n_{k})_{k\geq 0}$ and a limiting process $(X,\kappa,L)$ such that:
@@
Finally, the limit $(X,\kappa)$ satisfies the original reflected SDEL (4.49). The convergence holds in the Skorokhod topology, and the stochastic integrals converge by continuity. ∎
@@ 4.2 Markovian generalized doubly SDE
For all $(t,x)\in[0,T]\times\bar{G}$, let $(X^{t,x}_{s},\kappa^{t,x}_{s})_{s\geq 0}$ represents the solution to the reflected SDEL (4.49). Assume that the measurable functions $f,g,h$ and $H$ satisfy:
@@
As demonstrated in [12] (see Proposition 6.2), we will provide auxiliary estimations.
@@
There exists a non-negative constant $C$ such that for all $t,t^{\prime}\in[0,T]$ and $x,x^{\prime}\in\bar{G}$,
@@
Given $0,T]\times\bar{G}$ is bounded, $X^{t,x}_{s}\in\bar{G},\forall s\in[0,T]$ and the functions $f,g,h,H$ are continuous, it follows that there exists a constant $C$ such that
@@
$\mathbb{E}\left[\Phi_{T}^{\theta,\mu}\left|H(X^{t,x}_{T})\right|^{2}\right]\leq C\mathbb{E}\left[\Phi_{T}^{\theta,\mu}\left(1+\left|X^{t,x}_{T}\right|^{2}\right)\right]<\infty,$
@@
$\mathbb{E}\int_{0}^{T}\Phi_{s}^{\theta,\mu}\frac{\left|f(s,X^{t,x}_{s},0,0)\right|^{2}}{a^{2}_{s}}ds+\mathbb{E}\int_{0}^{T}\Phi_{s}^{\theta,\mu}\left|h(s,X^{t,x}_{s},0)\right|^{2}d\kappa^{t,x}_{s}+\mathbb{E}\int_{0}^{T}\Phi_{s}^{\theta,\mu}\left|g(s,X^{t,x}_{s},0)\right|^{2}ds<\infty.$
Consequently, as a result of Theorem 13 below, we obtain the following outcome.
@@
Then, employing the same reasoning as in [37] (see also [12]), we obtain the subsequent continuity result.
@@
Let $(Y_{s}^{t,x},Z_{s}^{t,x})$ be a solution to (4.65). The random field $(s,t,x)\rightarrow Y_{s}^{t,x},\ (s,t,x)\in[0,T]\times[0,T]\times\overline{\mathrm{G}}$ is almost surely continuous.
@@ 4.3 Stochastic viscosity solutions for SIPDE-NBC
We consider the stochastic IPDE with nonlinear Neumann boundary conditions of the form:
@@
$\mathcal{L}_{t}u(t,x)=b_{t}.\sigma(x).D_{x}u(t,x)+\int_{\mathbb{R}}\left[u(t,x+\sigma(x)e)-u(t,x)-D_{x}u(t,x)\sigma(x)e\right]F_{t}(de)$
$u^{1}_{k}(t,x)=\int_{\mathbb{R}}\left[u(t,x+\sigma(x)e)-u(t,x)-D_{x}u(t,x)e\right]p_{k}(e)F_{t}(de)$
@@
We now introduce the concept of stochastic viscosity solution for the SIPDE $(f,g,h)$. We are inspired by the work of Buckdahn and Ma [6], and we refer to their paper for a clear exposition on this topic. We use part of their nomenclature and follow the lines of their arguments to arrive at our major conclusion.
We shall utilize the stochastic flow $\hat{\chi}(t,x,y)\in\mathcal{C}(\mathbf{F}^{B},[0,T]\times\mathbb{R}^{l}\times\mathbb{R};\mathbb{R})$, which is the only solution to the SDE that looks like this in Stratonovich form:
@@
Under $(\mathcal{H}.3)$, the mapping $y\mapsto\hat{\chi}(t,x,y)$ constitutes a diffeomorphism for every $(t,x)$, $\mathbb{P}$-almost surely (see [42]). Let the $y$-inverse of $\hat{\chi}(t,x,y)$ be represented as $\tilde{\pi}(t,x,y)$. Consequently, given that $\tilde{\pi}(t,x,\hat{\chi}(t,x,y))=y$, it can be demonstrated (see [6] for instance)
@@
where the stochastic integrals must be read in the Stratonovich sense.
We will now introduce the process $\chi\in\mathcal{C}(\mathbf{F}^{B},[0,T]\times\mathbb{R}^{n}\times\mathbb{R};\mathbb{R})$ as the solution to the equation
@@
It is observed that, owing to the direction of the Itô integral, equation (4.56) must be seen as progressing from $T$ to $t$ (i.e., $y$ should be regarded as the beginning value). Consequently, the mapping $y\mapsto\chi(s,x,y)$ will exhibit same regularity characteristics as the mapping $y\mapsto\hat{\chi}(s,x,y)$ for all $(s,x)\in[t,T]\times\mathbb{R}^{n}$, $\mathbb{P}$-almost surely. Thus, if we indicate its $y$-inverse by $\pi$, we obtain
@@
    A random field $u\in\mathcal{C}\left(\mathbf{F}^{B},[0,T]\times\overline{\mathrm{G}};\mathbb{R}\right)$ is called a stochastic viscosity subsolution (resp. supersolution) of (4.55) if $u(T,x)\leq H(x)$ (resp. $\geq$), for all $x\in\overline{\mathrm{G}}$, and if for any stopping time $\tau\in\mathcal{M}_{0,T}^{B}$, any state variable $\xi\in L^{0}\left(\mathcal{F}_{\tau}^{B},G\right)$, and any random field $\varphi\in\mathcal{C}^{1,2}\left(\mathcal{F}_{\tau}^{B},[0,T]\times G;\mathbb{R}\right)$, with the property that for $\mathbb{P}^{2}$-almost all $\omega_{2}\in\{0<\tau<T\}$ the inequality
@@
    A random field $u\in\mathcal{C}\left(\mathbf{F}^{B},[0,T]\times\overline{\mathrm{G}};\mathbb{R}\right)$ is called a stochastic viscosity solution of (4.55) if it is both a stochastic viscosity subsolution and a supersolution.
@@
Note that if $f,h$ are deterministic and $g\equiv 0$, then Definition 20 coincides with the deterministic case (see, e.g., [12]).
Now, let us recall the concept of a random viscosity solution, which will serve as a connection between the stochastic viscosity solution and its deterministic equivalent.
@@
A random field $u\in\mathcal{C}\left(\mathbf{F}^{B},[0,T]\times\mathbb{R}^{l};\mathbb{R}\right)$ is called an $\omega$-wise viscosity solution if for $\mathbb{P}$-almost all $\omega\in\Omega,u(\omega,\cdot,\cdot)$ is a (deterministic) viscosity solution of the $\operatorname{SIPDE}(f,0,h)$.
To proceed, let us present the following lemma adapted from [27].
@@
$|\bar{h}(s,y)|\leq\tilde{a}_{s}(e^{2}\wedge|e|)\quad\text{a.s.},$
where $(\tilde{a}_{s})_{s\geq 0}$ is a positive predictable process with $\mathbb{E}\left[\int_{0}^{T}\tilde{a}_{s}^{2}ds\right]<\infty$. Then
$\sum_{t\leq s\leq T}\bar{h}(s,\Delta L_{s})=\sum_{i=1}^{d}\int_{t}^{T}\langle\bar{h}(s,\cdot),p_{i}\rangle_{L^{2}(\nu)}dH_{s}^{(i)}+\int_{t}^{T}\int_{\mathbb{R}}\bar{h}(s,e)F_{s}(de)ds,$
@@
$\langle\bar{h}(s,\cdot),p_{i}\rangle_{L^{2}(\nu)}=\int_{\mathbb{R}}\bar{h}(s,e)p_{i}(e)F_{s}(de)ds.$
@@
Recall that under assumption $(\mathcal{H}.3)$, the random field $\chi$ belongs to $C^{0,2,2}\left(\mathbf{F}^{B}\right$., $0,T]\times\mathbb{R}^{n}\times\mathbb{R};\mathbb{R}$) and hence, it is true for $\pi$. Subsequently, by using the transformation $\psi(t,x)=\chi(t,x,\varphi(t,x))$, we derive
$D_{x}\psi$ $=D_{x}\chi+D_{y}\chi D_{x}\varphi,$
@@
where all the derivatives of the random field $\pi(\cdot,\cdot,\cdot)$ are evaluated at $(t,x,\chi(t,x,y)$), whereas all derivatives of $\chi(\cdot,\cdot,\cdot)$ are evaluated at $(t,x,y)$.
@@
$\mathcal{L}_{t}\psi(t,x)$ $=b_{t}.\sigma(x).D_{x}\psi(t,x)+\int_{\mathbb{R}}\left[\psi(t,x+\sigma(x)e)-\psi(t,x)-D_{x}\psi(t,x)\sigma(x)e\right]F_{t}(de)$
@@
$+\int_{\mathbb{R}}\Big[\chi(t,x+\sigma(x)e,\varphi(t,x+\sigma(x)e))-\chi(t,x+\sigma(x)e,\varphi(t,x))-D_{y}\chi\cdot D_{x}\varphi\cdot\sigma(x)e\Big]F_{t}(de).$
@@
$+\int_{\mathbb{R}}\Bigl[\chi\bigl(t,x+\sigma(x)e,\,\varphi(t,x+\sigma(x)e)\bigr)-\chi-\bigl(D_{x}\chi+D_{y}\chi\cdot z\bigr)\sigma(x)e\Bigr]F_{t}(de)\Bigg]$ (4.60)
@@
$\tilde{\psi}^{1}_{k}(t,x)=\int_{\mathbb{R}}\left[\chi(t,x+\sigma e,\varphi(t,x+\sigma e))-\chi-(D_{x}\chi+D_{y}\chi D_{x}\varphi)\cdot e\right]p_{k}(e)F_{t}(de)$
@@
To establish part (b) in Definition 20, we notice that for every $(t,x)\in[0,T]\times\partial\mathrm{G}$, the subsequent series of equalities is valid:
@@
By integrating (4.61) and (4.63), we establish that the random field $v$ is a stochastic viscosity subsolution of $\operatorname{SIPDE}(\widetilde{f},0,\widetilde{h}), the$refore concluding the proof of Proposition 25. ∎
@@
Assume $\left(\mathcal{H}.{5}\right)$ holds. Let $\chi$ be the unique solution to $\operatorname{SDE}$ (4.56) and $\pi$ be the $y$-inverse of $\chi$. Then, there exists a constant $C>0$, depending only on the bound of $g$ and its partial derivatives, such that for $\zeta=\chi,\pi$, the following inequalities hold $\mathbb{P}$-almost surely for all $(t,x,y)\in[0,T]\times\mathbb{R}^{l}\times\mathbb{R}$:
@@
${V_{s}^{t,x}=\int_{\mathbb{R}}\Bigl[\chi\bigl(t,X_{s}^{t,x}+\sigma(X_{s}^{t,x})e,\,Y_{s}^{t,x}\bigr)-\chi-\bigl(D_{x}\chi+D_{y}\chi\cdot Z^{t,x}_{s}\bigr)\sigma(X_{s}^{t,x})e\Bigr]F_{s}(de)}$
Then, from Proposition 26, we obtain
$\left(\left(U_{s}^{t,x},V_{s}^{t,x}\right),(s,x)\in[0,T]\times\overline{\mathrm{G}}\right)\in\mathcal{S}^{2}_{\theta,\mu}\times\mathcal{H}^{2,\ell^{2}}_{\theta,\mu}$
@@
For each $(t,x)\in[0,T]\times\overline{\mathrm{G}}$, the process $\left(\left(U_{s}^{t,x},V_{s}^{t,x}\right),(s,x)\in[0,T]\times\overline{\mathrm{G}}\right)$ is the unique solution to the following generalized BSDE:
@@
where $\tilde{f}$ and $\tilde{h}$ are given by (4.3) and (4.62).
Here, the argument adheres to the same methodology as outlined in [4] (Theorem 3.4), together with Lemma 6. Therefore, we omit the proof.
@@
$u(\omega,t,x)=\chi(\omega,t,x,v(t,\omega,x)),\quad v(\omega,t,x)=\pi(\omega,t,x,u(t,\omega,x)).$
Subsequently, we obtain the following result, whose proof adheres to the same reasoning as shown in [4].
@@
The random field $u(t,x)=Y^{t,x}_{t}$ is a stochastic viscosity solution to the $\operatorname{SIPDE}(f,g,h)$.

### v3

@@ 4.1 Reflected SDE driven by inhomogeneous Lèvy process
Let $G$ be an open connected bounded domain of $\mathbb{R}^{l}$ $(l\geq 1)$. We suppose that $G$ is a smooth domain, which is such that for a function $\Psi\in C^{2}_{b}(\mathbb{R}^{l})$, $G$ and $\partial G$ are characterized by $G=\left\{x\in\mathbb{R}^{l},\Psi(x)>0\right\}$ and $\partial G=\left\{x\in\mathbb{R}^{l},\Psi(x)=0\right\},$ and for any $x\in\partial G$, $\nabla\Phi(x)$ is the unit normal vector pointing toward the interior of $G$. Furthermore, the interior sphere condition holds (see [37], pp.551), i.e. there exists $m>0$ such that for any $x\in\partial G,x^{\prime}\in\bar{G},$
@@
For every $(t,x)\in[0,T]\times\bar{G}$, we consider the following reflected stochastic differential equation driven by the inhomogeneous Lèvy process $L$:
@@
where $\sigma:\mathbb{R}^{l}\to\mathbb{R}^{l}$ is globally Lipschitz: $\exists C>0,\forall x,x^{\prime}\in\mathbb{R}^{l},\|\sigma(x)-\sigma(x^{\prime})\|\leq C|x-x^{\prime}|,$ and satisfies the linear growth condition: $\|\sigma(x)\|\leq C(1+|x|)$.
@@
Moreover, we have the following result of existence and uniqueness of the solution of (4.49). Here, the argument follows the same procedure as in [22]. For reader convenience we outline the main steps of the proof.
@@
The uniqueness result follows from Proposition 15. For the existence, we proceed as follows. To simplify the notation in the proof of the above proposition, we will denote $\mathfrak{R}=\mathfrak{R}^{t,x}$ for $\mathfrak{R}=X,\kappa$. For each $n\in\mathbb{N}$, define the truncated process:
$L_{t}^{n}=\int_{0}^{t}b_{s}ds+\int_{0}^{t}\int_{\frac{1}{n}<|z|<1}z\mu(ds,dz)-\int_{0}^{t}\int_{\frac{1}{n}<|z|<1}zF_{s}(dz)ds$ (4.50)
Since $L^{n}$ has finite variation, the Lipschitz and growth conditions hold on $\sigma$, and the domain $G$ satisfies the interior sphere condition (4.48), then from [22], for each $n$, there exists a unique solution $(X^{n},\kappa^{n})$ to the truncated problem:
$X_{s}^{n}=x+\int_{t}^{s}\sigma(X_{r-}^{n})dL_{r}^{n}+\int_{t}^{s}\nabla\Psi(X_{r}^{n})d\kappa_{r}^{n}$ (4.51)
Moreover, by Proposition 15 and Jakubowski’s compactness criterion for processes with right-continuous with left limits paths (see, e.g., [21]), there exists a subsequence $(n_{k})_{k\geq 0}$ and a limit process $(X,\kappa,L)$ such that:
@@
Finally, the limit $(X,\kappa)$ satisfies the original reflected SDE (4.49). The convergence holds in the Skorokhod topology, and the stochastic integrals converge by continuity. ∎
@@ 4.2 Markovian generalized doubly SDE
For all $(t,x)\in[0,T]\times\bar{G}$, let $(X^{t,x}_{s},\kappa^{t,x}_{s})_{s\geq 0}$ represents the solution to the reflected SDE (4.49). Assume that the measurable functions $f,g,h$ and $H$ satisfy:
@@
First, as in [12] (see Proposition 6.2) we shall give an auxiliary estimations.
@@
There exist a non-negative constant $C$ such that for all $t,t^{\prime}\in[0,T]$ and $x,x^{\prime}\in\bar{G}$,
@@
Since $0,T]\times\bar{G}$ is bounded, $X^{t,x}_{s}\in\bar{G},\forall s\in[0,T]$ and the functions $f,g,h,H$ are continuous, then there exists a constant $C$ such that
@@
$\mathbb{E}\left[\Phi_{T}^{\theta,\mu}\left|H(X^{t,x}_{T})\right|^{2}\right]\leq C\mathbb{E}\left[\Phi_{T}^{\theta,\mu}\left(1+\left|X^{t,x}_{T}\right|^{2}\right)\right]<\infty,$ (4.53)
@@
$\mathbb{E}\int_{0}^{T}\Phi_{s}^{\theta,\mu}\frac{\left|f(s,X^{t,x}_{s},0,0)\right|^{2}}{a^{2}_{s}}ds+\mathbb{E}\int_{0}^{T}\Phi_{s}^{\theta,\mu}\left|h(s,X^{t,x}_{s},0)\right|^{2}d\kappa^{t,x}_{s}+\mathbb{E}\int_{0}^{T}\Phi_{s}^{\theta,\mu}\left|g(s,X^{t,x}_{s},0)\right|^{2}ds<\infty.$ (4.54)
Therefore, as a consequence of Theorem 14 below we have the following result.
@@
Using the same arguments as in [36], we get the following continuity result.
@@ 4.3 Stochastic viscosity solutions for SIPDE-NBC
In this part, we introduce the notion of stochastic viscosity solutions to SIPDE with nonlinear Neumann boundary conditions. We consider the stochastic IPDE with nonlinear Neumann boundary condition of the form:
@@
$\mathcal{L}_{t}u(t,x)=b_{t}.\sigma(x).D_{x}u(t,x)+\int_{\mathbb{R}}\left[u(t,x+\sigma(x)z)-u(t,x)-D_{x}u(t,x)\sigma(x)z\right]F_{t}(dz)$
$u^{1}_{k}(t,x)=\int_{\mathbb{R}}\left[u(t,x+\sigma(x)z)-u(t,x)-D_{x}u(t,x)z\right]p_{k}(z)F_{t}(dz)$
@@
We now define the notion of stochastic viscosity solution for the SIPDE $(f,g,h)$. We are inspired by the work of Buckdahn and Ma [6] and we refer to their paper for a lucid discussion on this topic. We use some of their notation and follow the lines of their proofs to obtain our main result.
Indeed, we will use the stochastic flow $\hat{\chi}(t,x,y)\in C(\mathbb{F}^{B},[0,T]\times\mathbb{R}^{n}\times\mathbb{R})$, defined as the unique solution of the SDE which, in Stratonovich form, reads as follows:
@@
Under the assumption $(\mathcal{H}.3)$ the mapping $y\mapsto\hat{\chi}(t,x,y)$ defines a diffeomorphism for all $(t,x)$, $\mathbb{P}$-almost surely (see [41]). Denote the $y$-inverse of $\hat{\chi}(t,x,y)$ by $\tilde{\pi}(t,x,y)$. Then, since $\tilde{\pi}(t,x,\hat{\chi}(t,x,y))=y$, one can show that (cf. Buckdahn and Ma, 2001)
@@
where the stochastic integrals have to be interpreted in Stratonovich sense.
Now, let us introduce the process $\chi\in C(\mathbb{F}^{B},[0,T]\times\mathbb{R}^{n}\times\mathbb{R})$ as the solution to the equation
@@
We note that due to the direction of the Itô integral, equation (4.58) should be viewed as going from $T$ to $t$ (i.e., $y$ should be understood as the initial value). Then $y\mapsto\chi(s,x,y)$ will have the same regularity properties as those of $y\mapsto\hat{\chi}(s,x,y)$ for all $(s,x)\in[t,T]\times\mathbb{R}^{n}$, $\mathbb{P}$-almost surely. Hence if we denote by $\pi$ its $y$-inverse, we obtain
@@
    A random field $u\in\mathcal{C}\left(\mathbf{F}^{B},[0,T]\times\overline{\mathrm{G}}\right)$ is called a stochastic viscosity subsolution (resp. supersolution) of (4.57) if $u(T,x)\leq H(x)$ (resp. $\geq$), for all $x\in\overline{\mathrm{G}}$, and if for any stopping time $\tau\in\mathcal{M}_{0,T}^{B}$, any state variable $\xi\in L^{0}\left(\mathcal{F}_{\tau}^{B},G\right)$, and any random field $\varphi\in\mathcal{C}^{1,2}\left(\mathcal{F}_{\tau}^{B},[0,T]\times G\right)$, with the property that for $\mathbb{P}^{2}$-almost all $\omega_{2}\in\{0<\tau<T\}$ the inequality
@@
    A random field $u\in\mathcal{C}\left(\mathbf{F}^{B},[0,T]\times\overline{\mathrm{G}}\right)$ is called a stochastic viscosity solution of (4.57) if it is both a stochastic viscosity subsolution and a supersolution.
@@
Observe that if $f,h$ are deterministic and $g\equiv 0$ then Definition 21 coincides with the deterministic case (see e.g. [12]).
Now, let us recall a notion of random viscosity solution which will be a bridge linking the stochastic viscosity solution and its deterministic counterpart.
@@
A random field $u\in\mathcal{C}\left(\mathbf{F}^{B},[0,T]\times\mathbb{R}^{l}\right)$ is called an $\omega$-wise viscosity solution if for $\mathbb{P}$-almost all $\omega\in\Omega,u(\omega,\cdot,\cdot)$ is a (deterministic) viscosity solution of the $\operatorname{SIPDE}(f,0,h)$.
Let us introduce the following lemma adapted from [26].
@@
$|\bar{h}(s,y)|\leq a_{s}(y^{2}\wedge|y|)\quad\text{a.s.},$
where $(a_{s})_{s\geq 0}$ is a positive predictable process with $\mathbb{E}\left[\int_{0}^{T}a_{s}^{2}ds\right]<\infty$. Then
$\sum_{t\leq s\leq T}\bar{h}(s,\Delta L_{s})=\sum_{i=1}^{d}\int_{t}^{T}\langle\bar{h}(s,\cdot),p_{i}\rangle_{L^{2}(\nu)}dH_{s}^{(i)}+\int_{t}^{T}\int_{\mathbb{R}}\bar{h}(s,y)F_{s}(dy)ds,$
@@
$\langle\bar{h}(s,\cdot),p_{i}\rangle_{L^{2}(\nu)}=\int_{\mathbb{R}}\bar{h}(s,y)p_{i}(y)F_{s}(dy)ds.$
@@
Let us recall that under assumption $(\mathcal{H}.3)$ the random field $\chi$ belongs to $C^{0,2,2}\left(\mathbf{F}^{B}\right$., $0,T]\times\mathbb{R}^{n}\times\mathbb{R}$ ), and hence that the same is true for $\pi$. Then, considering the transformation $\psi(t,x)=\chi(t,x,\varphi(t,x))$, we obtain
$D_{x}\psi$ $=D_{x}\chi+D_{y}\chi D_{x}\varphi$
@@
where all the derivatives of the random field $\pi(\cdot,\cdot,\cdot)$ are evaluated at $(t,x,\chi(t,x,y)$), and all those of $\chi(\cdot,\cdot,\cdot)$ are evaluated at $(t,x,y)$.
@@
$\mathcal{L}_{t}\psi(t,x)$ $=b_{t}.\sigma(x).D_{x}\psi(t,x)+\int_{\mathbb{R}}\left[\psi(t,x+\sigma(x)z)-\psi(t,x)-D_{x}\psi(t,x)\sigma(x)z\right]F_{t}(dz)$
@@
$+\int_{\mathbb{R}}\Big[\chi(t,x+\sigma(x)z,\varphi(t,x+\sigma(x)z))-\chi(t,x+\sigma(x)z,\varphi(t,x))-D_{y}\chi\cdot D_{x}\varphi\cdot\sigma(x)z\Big]F_{t}(dz).$
@@
$+\int_{\mathbb{R}}\Bigl[\chi\bigl(t,x+\sigma(x)r,\,\varphi(t,x+\sigma(x)r)\bigr)-\chi-\bigl(D_{x}\chi+D_{y}\chi\cdot z\bigr)\sigma(x)r\Bigr]F_{t}(dr)\Bigg]$ (4.62)
@@
$\tilde{\psi}^{1}_{k}(t,x)=\int_{\mathbb{R}}\left[\chi(t,x+\sigma r,\varphi(t,x+\sigma r))-\chi-(D_{x}\chi+D_{y}\chi D_{x}\varphi)\cdot r\right]p_{k}(r)F_{t}(dr)$
@@
In order to establish part (b) in Definition 21 we notice that for all $(t,x)\in[0,T]\times\partial\mathrm{G}$, the following string of equalities holds:
@@
[…]
The main objective of this section is to show how a SIPDE with coefficients $(f,g,h)$ is related to equation (1.5) introduced in Section 1.
@@
Assume $\left(\mathcal{H}.{5}\right)$ holds. Let $\chi$ be the unique solution to $\operatorname{SDE}$ (4.58) and $\pi$ be the $y$-inverse of $\chi$. Then, there exists a constant $C>0$, depending only on the bound of $g$ and its partial derivatives, such that for $\zeta=\chi$, and $\zeta=\pi$, the following inequalities hold $\mathbb{P}$-almost surely for all $(t,x,y)\in[0,T]\times\mathbb{R}^{l}\times\mathbb{R}$:
@@
${V_{s}^{t,x}=\int_{\mathbb{R}}\Bigl[\chi\bigl(t,X_{s}^{t,x}+\sigma(X_{s}^{t,x})r,\,Y_{s}^{t,x}\bigr)-\chi-\bigl(D_{x}\chi+D_{y}\chi\cdot Z^{t,x}_{s}\bigr)\sigma(X_{s}^{t,x})r\Bigr]F_{s}(dr)}$
Then, from Proposition 4.5, we obtain
$\left(\left(U_{s}^{t,x},V_{s}^{t,x}\right),(s,x)\in[0,T]\times\overline{\mathrm{G}}\right)\in\mathcal{S}^{2}(F;[0,T];\mathbb{R})\times\mathcal{M}^{2}\left(F;[0,T];\mathbb{R}^{d}\right)$
@@
For each $(t,x)\in[0,T]\times\overline{\mathrm{G}}$, the process $\left(U_{s}^{t,x},V_{s}^{t,x},t\leq s\leq T\right)$ is the unique solution to the following generalized BSDE:
@@
where $\tilde{f}$ and $\tilde{h}$ are given by (24) and (26).
Here, the argument follows the same procedure as in [4] (Theorem 3.4). Therefore, we omit the proof.
@@
$u(t,\omega,x)=\eta(\omega,t,x,v(t,\omega,x)),\quad v(t,\omega,x)=\varepsilon(\omega,t,x,u(t,\omega,x))$ (40)
@@
The random field $u$ is a stochastic viscosity solution to the $\operatorname{SIPDE}(f,g,h)$.
The proof follows the same arguments in [4].